1.let x<0.find the vector n=<x,y,z> that is orthogonal to all three vectors u=<1,1,-2>,v=<-1,2,0> and w=<-1,0,1>.

2.find a unit vector that is orthogonal to both u=<0,-1,-1> and v=<1,0,-1>.


1
Expert's answer
2021-06-24T08:59:26-0400

(1) Let t be the fourth coordinate

For n to be orthogonal to the three vectors. Then,

"n=\begin{vmatrix}
 i & j &k &t\\
 1 & 1 &-2 &0\\
-1&2 &0 &0\\
-1& 0 &1 &0
\end{vmatrix}\\
n=t"

(2) The unit vector can be find by first getting the vector

"n_1=\begin{vmatrix} i & j &k \\ 0 & -1& -1\\ 1& 0& -1 \end{vmatrix}\\ n_1=i-j+k\\ \|n_1\|=\sqrt{1^2+(-1)^2+1^2}=\sqrt3\\ \widehat{n_1}=\frac{n_1}{\|n_1\|}\\ \widehat{n_1}=\frac{1}{\sqrt3}i-\frac{1}{\sqrt3}j+\frac{1}{\sqrt3}k\\"